OpenStudy (anonymous):

There are 500 tickets in a raffle. If you have 3 of these tickets and 5 prizes are to be given, what is the probability you will win at least one prize?

OpenStudy (anonymous):

The probability you win at least once is the same as one minus the probability you win nothing. That could be described as\[1-(\frac{497}{500})(\frac{496}{499})(\frac{495}{498})(\frac{494}{497})(\frac{493}{496})\]

OpenStudy (anonymous):

Each fraction is the probability you will lose that (first, second, etc.) drawing.

OpenStudy (anonymous):

Since they are independent events, we multiply. I leave the calculation to you.

OpenStudy (anonymous):

Thank you so much!!